OpenStudy (anonymous):

A die is rolled 20 times and the number of twos that come up is tallied. Find the probability of getting exactly four twos. A. 0.202 B. 0.075 C. 0.101 D. 0.083

OpenStudy (anonymous):

C

OpenStudy (anonymous):

\[\sqrt[\lim_{\lim_{\lim_{\lim_{\leftarrow \int\limits_{\sum_{\rightarrow}^{?}}^{?} \rightarrow ?} \rightarrow ?} \rightarrow ?} \rightarrow ?}]{?}\]

jimthompson5910 (jim_thompson5910):

Use the binomial distribution probability formula P(X = x) = (n ncr x)*(p)^(x)*(1-p)^(n-x) P(X = 4) = (20 ncr 4)*(0.1667)^(4)*(1-0.1667)^(20-4) P(X = 4) = (20 ncr 4)*(0.1667)^(4)*(0.8333)^(20-4) P(X = 4) = (4845)*(0.1667)^(4)*(0.8333)^16 P(X = 4) = (4845)*(0.000772)*(0.054053) P(X = 4) = 0.202177 So the probility is 0.202, which means that the answer is choice A

OpenStudy (anonymous):

Jim are you good at graphs by any chance?

OpenStudy (anonymous):

the answer is not c